# Difficulty-Aware Agent Orchestration in LLM-Powered Workflows

### v2

@@ Introduction
where $\mathbb{P}(\mathcal{G}\mid\mathcal{Q})$ is a distribution over query-specific workflows, $U(\cdot)$ and $C(\cdot)$ denote the utility (e.g., accuracy) and cost (e.g., token usage, latency) of executing workflow $\mathcal{G}$ on query $\mathcal{Q}$, respectively, and $\lambda$ is a trade-off coefficient that balances performance and efficiency.
Methodology
We first provide a roadmap of our method, then elaborate on each component in detail. Figure 1 presents the overall architecture of DAAO, which dynamically generates agentic workflows for input queries from various domains and difficulty levels. A controller network (Zoph and Le 2017) analyzes each query’s complexity and domain features, and dispatches it to three key modules: a query difficulty estimator $(\mathbb{N}{\theta_{d}})$, an agentic operator allocator $(\mathbb{N}{\theta_{p}})$, and an LLM router $(\mathbb{N}_{\theta_{m}})$. The estimator, based on a variational autoencoder (VAE) (Kingma and Welling 2014), encodes the query into a latent difficulty representation, guiding workflow depth and operator choice. The allocator selects suitable agentic operators using this representation and context, while the LLM router assigns appropriate language models by balancing reasoning needs and computational cost. This process yields a customized, multi-stage workflow for each query. After execution, output quality is evaluated, and feedback is used to jointly update the controller and modules, enabling continual improvement over time.
Difficulty-Aware Agent Orchestration
Given a query $\mathcal{Q}$, DAAO aims to dynamically construct a customized agentic workflow by analyzing the query’s complexity and domain characteristics, with the objective of generating a high-quality response:
$p(a|\mathcal{Q})=\int e(a|\mathcal{G})\mathbb{N}_{\theta}(\mathcal{G}|\mathcal{Q})d\mathcal{G},$ (6)
where $\mathbb{N}_{\theta}$ denotes the controller network parameterized by $\theta$, which takes $\mathcal{Q}$ as input. The function $e(\cdot|\cdot)$ denotes the conditional likelihood of obtaining the response $a$ by executing the workflow $\mathcal{G}$. The controller network $\mathbb{N}_{\theta}$ is composed of three sequential modules:
$\mathbb{N}_{\theta}=\mathbb{N}_{\theta_{d}}\circ\mathbb{N}_{\theta_{p}}\circ\mathbb{N}_{\theta_{m}},$ (7)
where $\mathbb{N}_{\theta_{d}}:\mathcal{Q}\rightarrow z$ is the query difficulty estimator, $\mathbb{N}_{\theta_{p}}:(\mathcal{Q},\mathbb{O},z)\rightarrow\{\mathcal{V}_{i}\}_{i=1}^{L}$ is the agentic operator allocator, and $\mathbb{N}_{\theta_{m}}:(\mathcal{Q},z,\{\mathcal{V}_{i}\}_{i=1}^{L})\rightarrow\{\mathcal{V}^{j}\}_{j=1}^{L}$ is the LLM router, which assigns specific language models to each selected operator. Here, the operator $\circ$ denotes function composition, applied in order from right to left.
Query Difficulty Estimator.
The module $\mathbb{N}_{\theta_{d}}:\mathcal{Q}\rightarrow z$ maps the input query $\mathcal{Q}$ to a latent difficulty representation $z\in\mathbb{R}^{k}$ using a variational autoencoder (VAE). Given a query embedding $x\in\mathbb{R}^{d}$, the encoder network parameterized by $\theta_{d}$ outputs the mean and log-variance of a Gaussian posterior:
$\mu,\log\sigma^{2}=f_{\text{enc}}(x),\quad z\sim\mathcal{N}(\mu,\sigma^{2}),$ (8)
where $f_{\text{enc}}$ denotes the encoder. The latent variable $z$ is then decoded into a scalar difficulty score:
$d=f_{\text{dec}}(z),\quad d\in[0,1].$ (9)
To train the VAE, we employ a difficulty-guided objective that encourages alignment between the predicted difficulty $d$ and the observed task outcome $y\in\{0,1\}$. The training loss is defined as:
$\mathcal{L}_{\text{difficulty}}=\|d-\tilde{d}\|_{2}^{2}+\lambda\cdot D_{\text{KL}}\left(q(z\mid x)\|p(z)\right),$ (10)
where $\tilde{d}$ is a pseudo-target adjusted based on the outcome:
$\tilde{d}=\text{clamp}\left(d+\gamma(1-2y),\;0,\;1\right).$ (11)
Here, $\gamma$ controls the adjustment magnitude, and $\lambda$ balances the prediction loss and KL regularization. This enables $\mathbb{N}_{\theta_{d}}$ to learn a meaningful prior over query difficulty, guiding downstream operator selection and model routing. In the KL term, $q(z\mid x)=\mathcal{N}(z;\mu,\sigma^{2})$ denotes the approximate posterior from the encoder, and $p(z)=\mathcal{N}(0,I)$ is the standard Gaussian prior. The KL divergence encourages $q(z\mid x)$ to stay close to $p(z)$, promoting regularity in the latent space.
Agentic Operator Allocator.
Before assigning agentic operators, we first determine the number of layers $L$ in the resulting workflow based on the predicted difficulty of the input query. Specifically, $L$ is computed as
$L=\left\lceil d\cdot\ell\right\rceil,$
where $\ell$ is a hyperparameter denoting the maximum allowed number of layers, and $d\in[0,1]$ is the predicted difficulty score from the previous module. This formulation allows the workflow complexity to be adaptively scaled with query difficulty.
We implement $\mathbb{N}_{\theta_{p}}$ as a sequential operator selection process:
$\mathbb{N}_{\theta_{p}}(\mathcal{G}\mid q,z,\mathbb{O})=\prod_{\ell=1}^{L}\pi_{\ell}(\mathcal{V}_{\ell}\mid q,z,\{\mathcal{V}_{h}\}_{h=1}^{\ell-1}),$ (12)
where $\mathcal{V}_{h}$ denotes the set of operators selected at layer $h$. The operator selection at layer $\ell$ is conditioned on the query $q$, latent difficulty $z$, and operator sets from preceding layers.
We implement the policy $\pi_{\ell}$ using a Mixture-of-Experts (MoE) architecture (Shazeer et al. 2017; Huang et al. 2024). For each layer $\ell$, the model selects a subset of operators $\mathcal{V}_{\ell}=\{\mathcal{O}_{\ell 1},\mathcal{O}_{\ell 2},\dots,\mathcal{O}_{\ell t}\}$ based on the current query and difficulty embedding:
$\pi_{\ell}:(q,z)\rightarrow\mathcal{V}_{\ell}.$ (13)
The number of selected operators $t$ is determined by thresholding the sorted activation scores:
$t=\underset{k\in\{1,\dots,|\mathbb{O}|\}}{\arg\min}\sum_{j<k}\mathbf{S}^{\downarrow}_{j}>P,$ (14)
where $\mathbf{S}\in\mathbb{R}^{|\mathbb{O}|}=[S_{1},\dots,S_{|\mathbb{O}|}]$ contains the activation scores of all candidate operators, and $\mathbf{S}^{\downarrow}$ denotes the same scores sorted in descending order. Operators are included sequentially until the accumulated score exceeds a predefined threshold $P$.
Each activation score $S_{i}$ is computed using a feed-forward network:
$S_{i}=\operatorname{FFN}\left(z\|\mathbf{v}(q)\|\sum_{\mathcal{O}\in\mathcal{V}_{1}}\mathbf{v}(\mathcal{O})\|\cdots\|\sum_{\mathcal{O}\in\mathcal{V}_{\ell-1}}\mathbf{v}(\mathcal{O})\right),$ (15)
where $\mathbf{v}(\cdot)$ is an embedding function implemented using lightweight models such as MiniLM (Wang et al. 2020) or Sentence-BERT (Reimers 2019), and $\|$ denotes vector concatenation.
@@
ComplexCoT (Fu et al. 2022) gemini-1.5-flash 80.74 86.01 48.28 80.12 71.80 73.39
@@
SC(CoT) (Wang et al. 2023a) gemini-1.5-flash 81.66 87.50 48.73 80.75 72.00 74.13
@@
ADAS (Hu, Lu, and Clune 2024) gemini-1.5-flash 79.68 86.00 45.89 80.69 68.00 72.05
@@
AFlow (Zhang et al. 2024b) gemini-1.5-flash 82.35 90.43 52.00 85.69 76.00 77.29
@@
MaAS (Zhang et al. 2025) gemini-1.5-flash 83.42 92.00 52.25 90.55 82.69 80.18
PromptLLM (Feng, Shen, and You 2024) LLM Pool 78.43 88.68 52.30 86.33 73.60 75.86
RouteLLM (Ong et al. 2024) LLM Pool 81.04 89.00 51.00 83.85 72.60 75.50
MasRouter (Yue et al. 2025) LLM Pool 84.25 92.00 52.42 90.62 84.00 80.66
@@
Table 1: Performance comparison across baseline prompting strategies, single-agent methods, autonomous agentic workflows, and LLM routing approaches. Bold numbers indicate the best performance, while underlined numbers denote the second-best. The LLM pool comprises both lightweight and high-capacity models to support diverse routing strategies.
LLM Router.
Inspired by (Ye et al. 2025; Barandoni et al. 2024), we recognize that different large language models (LLMs) offer varying strengths and limitations. Effective routing should exploit this heterogeneity rather than enforce uniformity. In workflows where all operators share the same LLM, even when enhanced with techniques such as chain-of-thought (CoT) or debate frameworks, the resulting reasoning process remains homogeneous. This often limits performance gains and increases computational cost. To address this, we adopt a heterogeneous workflow design where each operator is paired with the LLM best suited to its function. This approach preserves model diversity, improves efficiency, and enhances task performance through specialization.
We model the assignment of an LLM $\mathcal{M}_{i}$ to each selected operator $\mathcal{O}_{i}$, conditioned on the input query $\mathcal{Q}$ and its latent difficulty $z$. The overall routing objective is defined as:
$\mathbb{N}_{\theta_{m}}(\{\mathcal{M}_{i}\}_{i=1}^{k}\mid\mathcal{Q},z,\{\mathcal{O}_{i}\}_{i=1}^{k})=\prod_{i=1}^{k}\pi_{m}(\mathcal{M}_{i}\mid\mathcal{Q},z,\mathcal{O}_{i}),$ (16)
where $\pi_{m}(\mathcal{M}_{i}\mid\mathcal{Q},z,\mathcal{O}_{i})$ denotes the probability of assigning LLM $\mathcal{M}_{i}$ to operator $\mathcal{O}_{i}$.
Each selection probability is modeled using a temperature-scaled softmax over projected embeddings:
$\pi_{m}(\mathcal{M}_{i}\mid\mathcal{Q},z,\mathcal{O}_{i})=\frac{\exp\left(\frac{\langle H_{i}^{\text{comb}},e_{\mathcal{M}_{i}}\rangle}{\tau}\right)}{\sum_{j=1}^{N_{m}}\exp\left(\frac{\langle H_{i}^{\text{comb}},e_{\mathcal{M}_{j}}\rangle}{\tau}\right)},$ (17)
where $H_{i}^{\text{comb}}=\text{FFN}_{\text{comb}}\left([\text{FFN}_{q}([\mathcal{Q};W_{z}z]);\text{FFN}_{o}(\mathcal{O}_{i})]\right)\in\mathbb{R}^{d}$ is the combined contextual embedding of the query, difficulty, and operator. The term $e_{\mathcal{M}_{j}}=\text{FFN}_{m}(\mathcal{M}_{j})\in\mathbb{R}^{d}$ is the projected embedding of candidate LLM $\mathcal{M}_{j}$, and $\tau$ is a temperature parameter controlling the sharpness of the distribution. The dot product $\langle\cdot,\cdot\rangle$ measures cosine similarity after the embeddings are normalized.
The routing module is trained to maximize the likelihood of selecting the ground-truth LLMs:
$\mathcal{L}_{\text{llm}}=\sum_{i=1}^{k}\log\pi_{m}(\mathcal{M}_{i}\mid\mathcal{Q},z,\mathcal{O}_{i}).$ (18)
This formulation enables the system to route operators to diverse LLMs based on query difficulty and operator context, promoting specialized and adaptive reasoning across the workflow.
Experiments
Experiment Setup
@@
We evaluate DAAO on six public benchmarks covering three domains: (1) math reasoning, GSM8K (Cobbe et al. 2021) and MATH (Hendrycks et al. 2021b); (2) code generation, HumanEval (Chen et al. 2021) and MBPP (Austin et al. 2021)); tool use, GAIA (Mialon et al. 2023). Additionally, we include MMLU (Hendrycks et al. 2021a), a benchmark covering 57 academic subjects, to assess general knowledge and multitask language understanding. For the MATH benchmark, we follow (Hong et al. 2024) in selecting a harder subset (617 problems).
@@
We compare DAAO with three of agentic baselines: (1) single-agent approaches, including CoT (Wei et al. 2022), ComplexCoT (Fu et al. 2022), Self-Consistency (Wang et al. 2023a); (2) autonomous agentic workflows, including ADAS (Hu, Lu, and Clune 2024), AFlow (Zhang et al. 2024b) and MaAS (Zhang et al. 2025). (3) LLM routers, PromptLLM (Feng, Shen, and You 2024), RouteLLM (Ong et al. 2024) and MasRouter (Yue et al. 2025).
@@
We select LLM Pool with varying sizes and capacities, including gpt-4o-mini-0718 (OpenAI 2024), gemini-1.5-flash (Team et al. 2024), llama-3.1-70b (Dubey et al. 2024), Qwen-2-72b (Yang et al. 2024). LLMs are accessed via APIs, with the temperature set to 1.
@@
Building upon established methodologies in workflow automation (Saad-Falcon et al. 2024; Hu, Lu, and Clune 2024; Zhang et al. 2024b), we divide each dataset into training and test sets using a train:test ratio of 1:4. We initialize the feasible space of operator nodes with the following operators: CoT, LLM-Debate, Review, Ensemble, ReAct, Self-Consistency, Testing. We set the max number of layers as $L$ = 5, the cost penalty coefficient $\lambda$ as $\lambda\in\{1e-3,5e-3,1e-2\}$.
Performance Analysis
High-performing.
The experimental results in Table 1 demonstrate that DAAO effectively constructs high-performing agentic workflows. Compared to existing automated orchestration methods, DAAO achieves an average accuracy improvement of $2.83\%\sim 11.21\%$, and outperforms recent LLM routing methods by $2.60\%\sim 7.76\%$. On the MATH benchmark, DAAO attains a best-in-class score of 55.37%, surpassing the second-best method, MasRouter, by 2.95%. Across five datasets, DAAO consistently outperforms all baselines, highlighting its versatility and robustness.

### v3

@@ 3.1. Preliminary
where $\mathbb{P}(G|\mathcal{Q})$ is a distribution over query-specific workflows, $U(\cdot)$ and $C(\cdot)$ denote the utility (e.g., accuracy) and cost (e.g., token usage, latency) of executing workflow $G$ on query $\mathcal{Q}$, respectively, and $\lambda$ is a trade-off coefficient that balances performance and cost. The outer expectation is taken over the (empirical) data distribution of queries and answers, while the inner expectation marginalizes the stochastic workflow $G\sim P(\cdot\mid Q)$.
3.2. Difficulty-Aware Agent Orchestration
Given a query $\mathcal{Q}$, our DAAO builds a query-specific workflow by three difficulty-conditioned decisions produced by $N_{\theta_{L}}$ (workflow depths), $N_{\theta_{o}}$ (operator allocation), and $N_{\theta_{m}}$ (LLM selection).
Difficulty-conditioned decisions.
Each decision is a probability distribution whose logits are instantiated by the corresponding module and conditioned on the latent difficulty embedding $z$:
(6) $\underbrace{\pi^{(L)}(L\mid\mathcal{Q},z)}_{\text{workflow depth}(N_{\theta_{L}})},\quad\underbrace{\pi^{(O)}_{l}(O\mid\mathcal{Q},l,z)}_{\text{operator allocation}(N_{\theta_{o}})},\quad\underbrace{\pi^{(M)}(M\mid\mathcal{Q},O,z)}_{\text{model selection}(N_{\theta_{m}})}.$
Here $z$ is produced by our proposed difficulty estimator $N_{\theta_{d}}$, and a calibrated scalar $d$ is decoded from $z$ (more details in $3.$) In practice, $z$ parameterizes the logits of $N_{\theta_{L}}$$N_{\theta_{o}}$$N_{\theta_{m}}$, while $d$ serves as a scalar hardness prior used for thresholding and capacity scaling. Higher $d$ encourages larger-capacity workflows (e.g., more layers or activating more operators), whereas lower $d$ promotes conservative workflows.
$3.$ Query Difficulty Estimator
To make workflow generation difficulty-aware and balance performance vs. cost per query, we propose a difficulty estimator that guides the subsequent modules. The difficulty estimator $N_{\theta_{d}}$ maps the input query $\mathcal{Q}$ to a $k$-dimensional latent difficulty representation $z\in\mathbb{R}^{k}$ by using a variational autoencoder. To encode a given query $\mathcal{Q}$, a lightweight embedding layer $E_{\phi}$ produces a query embedding:
(7) $x=E_{\phi}(\mathcal{Q})\in\mathbb{R}^{h},$
where $h$ is the embedding dimension (e.g., h=384h{=}384). We then model a Gaussian posterior for the latent difficulty with diagonal covariance, which captures per-dimension uncertainty while remaining stable and efficient to train:
(8) $\mu(x)=W_{\mu}x+b_{\mu}\in\mathbb{R}^{k},\qquad\log\sigma^{2}(x)=W_{\sigma}x+b_{\sigma}\in\mathbb{R}^{k},$
with layer weights $W_{\mu},W_{\sigma}\in\mathbb{R}^{k\times h}$ and bias $b_{\mu},b_{\sigma}\in\mathbb{R}^{k}$, yielding the variational posterior (approximate posterior) over the latent difficulty $z$ given the query embedding $x$:
(9) $q(z\mid x)=\mathcal{N}(\mu(x),\,\mathrm{diag}(\sigma^{2}(x))),$
We sample $z$ using the reparameterization:
(10) $z\;=\;\mu(x)+\sigma(x)\odot\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,I_{k}),$
where $\odot$ denotes elementwise multiplication. All stochastic nodes are reparameterized; gradients are taken w.r.t. $\mu$, $\sigma$ via the pathwise estimator. The latent $z\in\mathbb{R}^{k}$ provides a rich difficulty embedding for downstream policies (depth, operator allocation, model selection), while we also decode an interpretable scalar difficulty $d\in(0,1)$ as task difficulty. Concretely, a one-hidden-layer MLP maps $z$ to $d$:
(11) $d=\operatorname{sigmoid}(W_{o}^{\top}\,\mathrm{ReLU}(W_{h}z+b_{h})+b_{o})\in(0,1).$
where $W_{h}\in\mathbb{R}^{m\times k}$, $b_{h}\in\mathbb{R}^{m\times 1}$, $W_{o}\in\mathbb{R}^{1\times m}$ and $b_{o}\in\mathbb{R}$ are the weights and biases of the hidden and output layers, respectively.
We train our difficulty estimator $N_{\theta_{d}}$ with a difficulty-guided objective that aligns the decoded difficulty $d$ with the observed outcome $y\in\{0,1\}$ (solved $y=1$, not solved $y=0$), while regularizing the latent embedding $z$:
(12) $\mathcal{L}_{\text{diff}}=\mathcal{L}_{\text{cal}}(d,y)+\lambda\,D_{\mathrm{KL}}\!\big(q(z|x)\|p(z)\big),\quad p(z)=\mathcal{N}(0,I),$
where, the KL term keeps the approximate posterior $q(z|x)$ close to the standard normal prior $p(z)$, stabilizing the latent space. We adopt a binary cross-entropy (BCE) as the difficulty-calibration term. Since a higher $d$ denotes a harder query (thus lower success probability), we calibrate the predicted success probability as $\hat{p}_{\text{succ}}=1-d$ and set:
(13) $\mathcal{L}_{\text{cal}}(d,y)=\mathrm{BCE}\big(\hat{p}_{\text{succ}},\,y\big)=-y\log(1-d)-(1-y)\log d.$
The coefficient $\lambda>0$ balances calibration and regularization.
3.4. Agentic Operator Allocator.
Given a query $\mathcal{Q}$, this module constructs a directed acyclic workflow $G=\{\mathcal{V},\mathcal{E}\}$ from a candidate operator pool $\mathbb{O}\subseteq\mathbb{M}\times\mathbb{S}$, targeting difficulty-aware orchestration that balances performance and cost.
Depth adaptation.
Let $L_{max}$ denote the maximum depth. We set
(14) $L=\max\{1,\,\lceil d\cdot L_{max}\rceil\},$
so that easier queries yield shallower graphs while harder ones trigger deeper chains.
Layer-wise MoE selection.
We select operators layer by layer using a lightweight MoE gate and factorize the joint selection over layers as an autoregressive process:
(15) $\mathcal{N}_{\theta_{o}}\!\big(G\mid\mathcal{Q},z,O\big)=\prod_{l=1}^{L}\,\pi_{l}^{(O)}\!\big(V_{l}\,\big|\,\mathcal{Q},z,V_{<l}\big),\quad V_{<l}:=\{V_{1},\dots,V_{l-1}\}.$
Here, $V_{<l}:=\{V_{1},\dots,V_{l-1}\}$ denotes the history (all previously chosen operator sets), and $\pi_{l}(\cdot)$ is the layer-$l$ Mixture-of-Experts (MoE) policy (Shazeer et al., 2017; Huang et al., 2024) that outputs a subset $V_{l}$. This factorization means “choose the current layer conditioned on all previous choices $V_{<l}$, enabling difficulty-aware $z$, query context-dependent $Q$ routing.”
Scoring-to-selection with adaptive width.
We instantiate $\pi_{l}^{(O)}$ so that each layer’s width (number of operators) adapts to the available evidence. First, we translate the query difficulty context and the history into scalar compatibilities for all candidates:
(16) $S_{i}\;=\;\operatorname{FFN}\!\Big(z\,\|\,v(\mathcal{Q})\,\|\,\sum_{O\in V_{1}}\!v(O)\,\|\,\cdots\,\|\,\sum_{O\in V_{l-1}}\!v(O)\Big),\quad O_{i}\in\mathbb{O},$
where $v(\cdot)$ is a lightweight embedding (e.g MiniLM (Wang et al., 2020) or Sentence-BERT (Reimers, 2019)) and $\|$ denotes concatenation. Larger $S_{i}$ means $i$-th operator $O_{i}$ is more compatible with the current layer, given the accumulated context.
We then convert scores into a subset using a cumulative-threshold decoder. Let $\mathbb{S}=[S_{1},\dots,S_{|\mathbb{O}|}]$, and $S_{(1)}\geq S_{(2)}\geq\cdots\geq S_{(|\mathbb{O}|)}$ be the scores in descending order. Using a preset threshold $\tau$, we determine the number of operators as:
(17) $t\;=\;\min\Big\{\,r\in\{1,\dots,|\mathbb{O}|\}\;:\;\sum_{i=1}^{r}S_{(i)}\;>\;\tau\,\Big\}.$
where $r$ is the prefix length (the number of top-ranked operators considered). Operators are activated in descending score order ($S_{(1)}\geq S_{(2)}\geq\cdots$) and the activation proceeds sequentially until the cumulative evidence exceeds $\tau$. This rule yields an adaptive layer width—higher aggregate confidence activates more operators; lower confidence activates fewer—while enforcing a budget-like constraint via $\tau$. Together, (16)–(17) provide a concrete instantiation of the per-layer policy.
3.5. LLM Router.
Inspired by (Ye et al., 2025; Barandoni et al., 2024), we leverage model heterogeneity rather than enforcing a single-LLM workflow. After operator selection, each chosen operator $O_{(i)}$ (for $i=1,\dots,t$) is paired with an LLM from a candidate set. We model the per-operator routing as
(18) $\mathbb{N}_{\theta_{m}}\!\big(\{M_{(i)}\}_{i=1}^{t}\,\big|\,\mathcal{Q},z,\{O_{(i)}\}_{i=1}^{t}\big)\;=\;\prod_{i=1}^{t}\pi^{(M)}\!\big(M_{(i)}\,\big|\,\mathcal{Q},z,O_{(i)}\big),$
where $\pi^{(M)}(\cdot\,|\,\mathcal{Q},z,O_{(i)})$ is the layer-agnostic LLM policy for the $i$-th selected operator. For each selected operator, we define a temperature-scaled softmax over LLM candidates indexed by $m\in\{1,\dots,N_{M}\}$:
(19) $\pi^{(M)}\!\big(M_{(i)}=M_{m}\,\big|\,\mathcal{Q},z,O_{(i)}\big)\;=\;\frac{\exp\!\left(\langle\hat{h}_{(i)},\,\hat{e}_{m}\rangle/T\right)}{\sum_{u=1}^{N_{M}}\exp\!\left(\langle\hat{h}_{(i)},\,\hat{e}_{u}\rangle/T\right)}\,.$
where $h_{(i)}\;=\;\mathrm{FFN}_{\mathrm{comb}}\!\big(\mathrm{FFN}_{q}(\mathcal{Q})\,\|\,W_{z}z\,\|\,\mathrm{FFN}_{o}(O_{(i)})\big)\in\mathbb{R}^{d}$, is the combined contextual embedding of the query, difficulty, and operator. $e_{m}\;=\;\mathrm{FFN}_{m}(M_{m})\in\mathbb{R}^{d}$ is the projected embedding of candidate LLM $M_{m}$ and $T$ is the temperature parameter controlling the sharpness of the distribution. The dot product $\langle\cdot,\cdot\rangle$ measures cosine similarity after the embeddings are normalized.
This routeing policy enables the system to route operators to diverse LLMs based on query difficulty and operator context, promoting specialized and adaptive reasoning across the workflow.
Table 1. Performance comparison across baseline prompting strategies, single-agent methods, autonomous agentic workflows, and LLM routing approaches. Bold numbers indicate the best performance, while underlined numbers denote the second-best. The LLM pool comprises both lightweight and high-capacity models to support diverse routing strategies.
@@ 3.2. Difficulty-Aware Agent Orchestration
ComplexCoT (Fu et al., 2022) gemini-1.5-flash 80.74 86.01 48.28 80.12 71.80 73.39
@@
SC(CoT) (Wang et al., 2023b) gemini-1.5-flash 81.66 87.50 48.73 80.75 72.00 74.13
@@
ADAS (Hu et al., 2024) gemini-1.5-flash 79.68 86.00 45.89 80.69 68.00 72.05
@@
AFlow (Zhang et al., 2024a) gemini-1.5-flash 82.35 90.43 52.00 85.69 76.00 77.29
@@
MaAS (Zhang et al., 2025) gemini-1.5-flash 83.42 92.00 52.25 90.55 82.69 80.18
PromptLLM (Feng et al., 2024) LLM Pool 78.43 88.68 52.30 86.33 73.60 75.86
RouteLLM (Ong et al., 2024) LLM Pool 81.04 89.00 51.00 83.85 72.60 75.50
MasRouter (Yue et al., 2025) LLM Pool 84.25 92.00 52.42 90.62 84.00 80.66
@@
4. Experiments
4.1. Experiment Setup
@@
We evaluate DAAO on six public benchmarks covering three domains: (1) math reasoning, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b); (2) code generation, HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021)); tool use, GAIA (Mialon et al., 2023). Additionally, we include MMLU (Hendrycks et al., 2021a), a benchmark covering 57 academic subjects, to assess general knowledge and multitask language understanding. For the MATH benchmark, we follow (Hong et al., 2024) in selecting a harder subset (617 problems). The dataset metric are in Appendix C.
@@
We compare DAAO with three of agentic baselines: (1) single-agent approaches, including CoT (Wei et al., 2022), ComplexCoT (Fu et al., 2022), Self-Consistency (Wang et al., 2023b); (2) autonomous agentic workflows, including ADAS (Hu et al., 2024), AFlow (Zhang et al., 2024a) and MaAS (Zhang et al., 2025). (3) LLM routers, PromptLLM (Feng et al., 2024), RouteLLM (Ong et al., 2024) and MasRouter (Yue et al., 2025).
@@
We select LLM Pool with varying sizes and capacities, including gpt-4o-mini-0718 (OpenAI, 2024), gemini-1.5-flash (Team et al., 2024), llama-3.1-70b (Dubey et al., 2024), Qwen-2-72b (Yang et al., 2024). LLMs are accessed via APIs, with the temperature set to 1. We selected gpt-4o-mini-0718 (OpenAI, 2024) and gemini-1.5-flash (Team et al., 2024), which performed well in Vanilla, as the models for other baselines.
@@
Building upon established methodologies in workflow automation (Saad-Falcon et al., 2024; Hu et al., 2024; Zhang et al., 2024a), we divide each dataset into training and test sets using a train:test ratio of 1:4. We initialize the feasible space of operator nodes with the following operators: CoT, LLM-Debate, Review, Ensemble, ReAct, Self-Consistency, Testing. Detailed instructions are in Appendix B.1. We set the max number of layers as $L_{max}$ = 5, the cost penalty coefficient $\lambda$ as $\lambda\in\{1e-3,5e-3,1e-2\}$, the sampling times $K$ = 4 and threshold $\tau$ = 0.3. To ensure robust results, we conducted each experiment three times and reported the average performance.
4.2. Performance Analysis
4.2.1. High-performing.
The experimental results in Table 1 demonstrate that DAAO effectively constructs high-performing agentic workflows. Compared to existing automated orchestration methods, DAAO achieves an average accuracy improvement of $3.5\%\sim 15.2\%$, and outperforms recent LLM routing methods by $3.2\%\sim 10.2\%$. On the MATH benchmark, DAAO attains a best-in-class score of 55.37%, surpassing the second-best method, MasRouter, by 2.95%. Across five datasets, DAAO consistently outperforms all baselines, highlighting its versatility and robustness.